Find a formula for the inverse of the function, if the function is one-to-one. If not, state the function is not one-to-one for
F(x)=x^3-1

Answer by CubeyThePenguin(3113)

Solve for x.

y = x^3 - 1
y + 1 = x^3
cbrt(y + 1) = x

f^(-1)(x) = cbrt(x + 1)